### v1

# Sustainable Exploitation Equilibria for Dynamic Games

Abstract
We introduce the Sustainable Exploitation Equilibrium (SEE), a refinement of Markov Perfect Equilibrium (MPE) for dynamic games with an exploiter–exploitee structure. SEE imposes two additional discipline conditions: (i) viability, requiring state trajectories to remain inside a sustainability set; and (ii) renegotiation-proofness with exploiter-optimal selection, to retain only those viable equilibria that are immune to Pareto-improving renegotiations, with ties resolved in favor of the exploiter. In our base formulation the exploitee cannot exit the relationship (no outside option), but retains a strategic effort margin that affects dynamics and payoffs. We establish existence under appropriate conditions and illustrate SEE in a hegemon-client model of foreign politics, where tribute demands trade off against the client’s governance effort.
Keywords: Markov perfect equilibrium, equilibrium refinements, viability, renegotiation-proofness, sustainability
1 Introduction
Exploitation arises when one agent extracts value from another whose continued functioning is essential for the relationship itself. Standard equilibrium concepts such as Markov Perfect Equilibrium (MPE) may admit outcomes in which the exploiter extracts too aggressively, driving the state outside the range in which the exploitee can survive. To address this, the paper introduces the Sustainable Exploitation Equilibrium (SEE), a refinement of stationary MPE that selects only those equilibria that (i) keep the state within a viability set and (ii) are immune to Pareto-improving renegotiations, with exploiter-optimal selection. The contribution is to formalize a simple and tractable equilibrium notion suited to persistent asymmetric relationships in which collapse is feasible but avoided in equilibrium. SEE is motivated by settings where one party cannot exit but remains responsive, such as in hegemon–client dynamics, as the exploitee retains a strategic effort margin that shapes both payoffs and dynamics. In this way, SEE formalizes the idea of “sustainable exploitation”: the exploiter extracts as much as possible while ensuring the long-run survival and productivity of the exploitee.
2 Related Literature
Our work connects three strands of existing literature.
Equilibrium refinements.
SEE builds on the refinement tradition in dynamic games. (MaskinTirole2001a) develop the theory of Markov perfect equilibrium; (Bernheim1987) introduce coalition-proof Nash equilibrium; (Selten1975) trembling-hand perfection; and (FarrellMaskin1989) renegotiation-proofness in repeated games.
Viability and dynamic control.
The viability constraint draws from (Aubin1991) Viability Theory and constrained Markov decision processes (Altman1999), ensuring state trajectories remain within sustainability sets. This connects to the resource economics literature, e.g. (Dasgupta1979), (Chichilnisky1997), and (Dockner2000), where long-run viability is central to dynamic exploitation of resources.
Political economy and relational contracts.
SEE also relates to political economy models of extraction and state capacity ((AcemogluRobinson2006); (BesleyPersson2011)) and to relational contracting without full commitment ((ThomasWorrall1988); (Levin2003)). In these contexts, weaker parties cannot exit but retain effort margins that shape dynamics. SEE formalizes the logic of “restrained exploitation” within a Markovian framework.
3 Model and Definition
Environment and timing.
Time is discrete, $t=0,1,2,\dots$. The state $s_{t}\in S$ evolves on a compact metric space $S$. In each period:
  1. 1.
    The exploiter $X$ observes $s_{t}$ and chooses extraction $x_{t}\in X(s_{t})$.
  2. 2.
    The exploitee $E$ observes $(s_{t},x_{t})$ and chooses effort $e_{t}\in E(s_{t})$.
  3. 3.
    The state transitions to $s_{t+1}\sim Q(\cdot\mid s_{t},x_{t},e_{t})$ (deterministic dynamics are $s_{t+1}=f(s_{t},x_{t},e_{t})$).
Per-period payoffs are $u^{X}(s_{t},x_{t},e_{t})$ and $u^{E}(s_{t},x_{t},e_{t})$, bounded and continuous. Both players discount with $\delta\in(0,1)$.
Roles of the players
Definition 1 (Exploiter).
The exploiter $X$ directly extracts value via $x\in X(s)$, appropriating resources and typically reducing the future state. $X$’s stage payoff $u^{X}(s,x,e)$ is weakly increasing in $x$ for fixed $(s,e)$. A Markov policy for $X$ is $\sigma^{X}:S\to X(s)$.
Definition 2 (Exploitee (no exit, effort choice)).
The exploitee $E$ cannot exit the relationship (no outside option) but chooses effort $e\in E(s)$ after observing $(s,x)$. Effort affects contemporaneous payoffs and/or the law of motion. A Markov policy for $E$ is $\sigma^{E}:S\times X(s)\to E(s)$ (second-mover response).
Equilibrium notion (Markov–Stackelberg Equilibrium).
A Markov–Stackelberg Equilibrium (MSE) consists of policies $(\sigma^{X},\sigma^{E})$ such that: (i) for each $(s,x)$, $\sigma^{E}(s,x)$ is a best response to $\sigma^{X}$ given follower timing; (ii) $\sigma^{X}(s)$ anticipates $\sigma^{E}$ and is a best response given $E$’s continuation behavior. Values $(W^{X},W^{E})$ solve the associated Bellman system.
Definition 3 (Sustainable Exploitation Equilibrium (SEE)).
Let $V\subseteq S$ be a nonempty compact viability set that is invariant under admissible state transitions. Let $\mathcal{E}$ denote the set of all stationary Markov Perfect Equilibria, and let $\mathcal{E}_{V}\subseteq\mathcal{E}$ denote the subset whose induced state process remains in $V$. We assume $\mathcal{E}_{V}\neq\emptyset$.
A Sustainable Exploitation Equilibrium is any stationary MPE $\sigma^{*}\in\mathcal{E}$ satisfying:
  1. (i)
    Viability: The state induced by $\sigma^{*}$ remains in the viability set $V$ for all $s\in V$ (i.e., $\sigma^{*}\in\mathcal{E}_{V}$).
  2. (ii)
    Renegotiation-proofness: At any state $s\in V$, there is no alternative viable profile $\tilde{\sigma}\in\mathcal{E}_{V}$ such that:
    $W_{E}(\tilde{\sigma};s)\geq W_{E}(\sigma^{*};s)\quad\text{and}\quad W_{X}(\tilde{\sigma};s)\geq W_{X}(\sigma^{*};s),$
    with at least one inequality strict.
  3. (iii)
    Exploiter-optimality: Among all renegotiation-proof elements of $\mathcal{E}_{V}$, $\sigma^{*}$ maximizes the exploiter’s value $W_{X}(\cdot;s)$ at the state $s$ (with ties broken lexicographically).
Any profile satisfying (i)–(iii) is a Sustainable Exploitation Equilibrium.
4 Relation to Nash and Markov Perfect Equilibrium
We position the Sustainable Exploitation Equilibrium (SEE) within the standard equilibrium hierarchy.
Nash equilibrium (NE).
A strategy profile $\sigma=(\sigma_{X},\sigma_{E})$ is a Nash equilibrium if no player can profitably deviate unilaterally. NE is broad and may rely on history-dependent strategies.
Subgame perfect equilibrium (SPE).
A refinement of NE, SPE requires sequential rationality in every subgame. This eliminates non-credible threats. Clearly,
$\text{SPE}\subseteq\text{NE}.$
Markov perfect equilibrium (MPE).
A further refinement of SPE for dynamic games with a payoff-relevant state $s\in S$. In MPE, strategies are Markovian: $\sigma_{i}(s)$ depends only on the current state, not the full history. For two-player sequential-move environments, the relevant benchmark is a Markov–Stackelberg equilibrium (MSE): the follower chooses $\sigma_{E}(s,x)$ as a best response given the state and leader’s action, while the leader anticipates this response in selecting $\sigma_{X}(s)$. Thus,
$\text{MPE (MSE)}\subseteq\text{SPE}\subseteq\text{NE}.$
Sustainable Exploitation Equilibrium (SEE).
SEE is defined as a viable, renegotiation-proof MPE with exploiter-optimal selection. Formally, a stationary Markov profile $(\sigma_{X},\sigma_{E})$ is an SEE if the aforementioned conditions of Definition 3 are satisfied.
It follows immediately that
$\text{SEE}\subseteq\text{MPE}\subseteq\text{SPE}\subseteq\text{NE}.$
5 Existence
Theorem 1 (Existence of SEE).
Suppose the state space $S$ and action sets $X(s),E(s)$ are compact; payoffs are bounded and continuous; and the transition kernel $Q$ satisfies the Feller property. Let $V\subseteq S$ be a nonempty compact viability set that is invariant under admissible state transitions. Assume that there exists at least one stationary Markov Perfect Equilibrium whose induced state process remains in $V$. Then:
(i) the set of viable stationary Markov Perfect Equilibria is nonempty and compact;
(ii) the subset of renegotiation-proof viable equilibria is nonempty; and
(iii) an exploiter-optimal selection exists.
Hence a Sustainable Exploitation Equilibrium (SEE) exists.
Sketch.
(1) Existence of stationary MPE. Under compact state and action spaces, bounded and continuous payoffs, and Feller continuity of the transition kernel, the stationary best-response correspondence is nonempty, convex-valued, and upper hemicontinuous. By Glicksberg’s fixed point theorem, there exists at least one stationary Markov Perfect Equilibrium (Maskin and Tirole, 2001).
(2) Viability. Let $\mathcal{E}$ denote the set of stationary MPE and define $\mathcal{E}_{V}\subseteq\mathcal{E}$ as the subset whose induced state process remains in the invariant compact set $V$. By assumption, $\mathcal{E}_{V}$ is nonempty. Compactness of $S$ and continuity of payoffs and transitions imply that the associated set of equilibrium value functions for $\mathcal{E}_{V}$ is compact in the product topology.
(3) Renegotiation-proofness. The set of feasible continuation values generated by $\mathcal{E}_{V}$ is compact. Therefore there exists at least one equilibrium in $\mathcal{E}_{V}$ that is not strictly Pareto-dominated by any other viable equilibrium; such an equilibrium is renegotiation-proof.
(4) Selection. The exploiter’s value function is continuous on the nonempty compact set of renegotiation-proof viable equilibria and hence attains a maximum. Any maximizer is a Sustainable Exploitation Equilibrium. ∎
6 Example: Foreign Politics (Hegemon–Client SEE)
6.1 Primitives and timing
Two players: a hegemon $H$ (the exploiter) and a client state $C$ (the exploitee). The state $s_{t}\in S\subset\mathbb{R}_{+}$ represents client political capacity/stability (e.g., tax capacity, administrative control, consent). Time is discrete $t=0,1,2,\dots$.
Within each period:
  1. 1.
    $H$ observes $s_{t}$ and chooses an extraction/pressure level $x_{t}\in X(s_{t})$ (e.g., tribute, resource concessions, policy compliance).
  2. 2.
    $C$ observes $(s_{t},x_{t})$ and chooses effort $e_{t}\in E(s_{t})$ (e.g., governance effort, compliance/implementation, regime maintenance).
  3. 3.
    The state updates to $s_{t+1}=f(s_{t},e_{t})-h(x_{t})$, where $f$ is increasing in $s$ and $e$ and $h$ is increasing in $x$; or more generally $s_{t+1}\sim Q(\cdot\mid s_{t},x_{t},e_{t})$.
Both discount with factor $\delta\in(0,1)$. The client cannot exit the relationship (no outside option) but has strategic effort. A viability requirement $s_{t}\geq s_{\min}$ avoids regime collapse or state failure.
6.2 Payoffs and feasibility
Per-period payoffs are continuous and bounded:
$u^{H}(s,x,e)\;=\;\pi(x)\;-\;k(s,e),\qquad u^{C}(s,x,e)\;=\;b(s)\;-\;\phi(e)\;-\;d(x),$
with $\pi^{\prime}(x)>0$, $k\geq 0$ (costs to monitor/coerce or instability spillovers), $b^{\prime}(s)\geq 0$, $\phi$ convex with $\phi^{\prime}(0)=0$, and $d^{\prime}(x)\geq 0$. Feasibility (viability) imposes
$x\;\leq\;f(s,e)-s_{\min}.$
6.3 Equilibrium notion and SEE
We adopt a Markov–Stackelberg Equilibrium (MSE): the follower $C$ best responds with $\sigma^{C}(s,x)$, and the leader $H$ best responds with $\sigma^{H}(s)$ anticipating $\sigma^{C}$. A Sustainable Exploitation Equilibrium (SEE) is a viable, renegotiation-proof MSE with exploiter-optimal selection (as in Definition 1).
6.4 Follower’s best response
Given $(s,x)$ and anticipating future play, $C$ solves
$\max_{e\in E(s)}\;\Big\{b(s)-\phi(e)-d(x)+\delta W^{C}\!\big(f(s,e)-h(x)\big)\Big\}.$
An interior first-order condition is
$\phi^{\prime}(e^{*})\;+\;\delta\,W^{C}_{s}\!\big(s^{\prime}\big)\,f_{e}(s,e^{*})\;=\;0,\qquad s^{\prime}\equiv f(s,e^{*})-h(x).$
Under standard concavity/regularity, the implicit function theorem yields $\partial e^{*}/\partial x$.
6.5 Leader’s SEE problem
Anticipating $e^{*}(s,x)$, the hegemon solves
$V^{H}(s)\;=\;\max_{x\in X(s)}\;\Big\{\pi(x)-k(s,e^{*}(s,x))+\delta\,V^{H}\!\big(f(s,e^{*}(s,x))-h(x)\big)\Big\}$
subject to viability $x\leq f(s,e^{*}(s,x))-s_{\min}$. Let $\mu\geq 0$ be the multiplier on viability. The Lagrangian is
$\mathcal{L}(x;s,\mu)\;=\;\pi(x)-k(s,e^{*})+\delta V^{H}(s^{\prime})\;+\;\mu\big(f(s,e^{*})-h(x)-s_{\min}\big),$
where $e^{*}\equiv e^{*}(s,x)$ and $s^{\prime}\equiv f(s,e^{*})-h(x)$.
The (interior) first-order condition with envelope terms from $e^{*}(s,x)$ is
$\pi^{\prime}(x)\;-\;\delta V^{H}_{s}(s^{\prime})\,\big(h^{\prime}(x)-f_{e}(s,e^{*})\,\partial e^{*}/\partial x\big)\;-\;\mu\,\big(h^{\prime}(x)-f_{e}(s,e^{*})\,\partial e^{*}/\partial x\big)\;=\;0,$
with complementary slackness $\mu\big(f(s,e^{*})-h(x)-s_{\min}\big)=0$.
Interpretation.
The hegemon equates marginal extraction benefits $\pi^{\prime}(x)$ to the discounted shadow cost of lost future capacity $h^{\prime}(x)$ net of the induced effort response $f_{e}\partial e^{*}/\partial x$. When viability binds, the shadow price $\mu>0$ further restrains extraction.
6.6 Stationary SEE and regimes
A stationary SEE consists of $(x^{*}(s),e^{*}(s,x))$ and an invariant distribution supported on $V=\{s\geq s_{\min}\}$. Two regimes arise:
  • Interior (viability slack, $\mu=0$). If the unconstrained fixed point $s^{*}$ of $s^{\prime}=f(s,e^{*}(s,x^{*}))-h(x^{*})$ satisfies $s^{*}>s_{\min}$, the FOCs imply a generalized “golden rule” where effective regeneration $f_{s}+f_{e}\,\partial e^{*}/\partial s$ balances discounting (details depend on primitives).
  • Boundary (viability binds, $\mu>0$). If the unconstrained $s^{*}$ would fall below $s_{\min}$, SEE selects $s^{*}=s_{\min}$ and $x^{*}=f(s_{\min},e^{*}(s_{\min},x^{*}))-s_{\min}$, with $\mu$ set to satisfy the FOC.
Additional Remark on Outside Option Expansions
One might alternatively assume that the exploitee $E$ has an outside option $U^{out}_{E}(s)$, in which case viability must be supplemented by an individual rationality (IR) constraint: $W^{E}(s)\geq U^{out}_{E}(s)$ for all $s\in V$. This yields an even tighter refinement, as it restricts feasible equilibria to those in which the exploitee weakly prefers participation over exit. We deliberately exclude outside options in the baseline SEE definition, since the core notion of exploitation is precisely that the exploitee cannot walk away. In other settings, the SEE framework can be adapted straightforwardly by adding the IR condition, producing a refinement that nests our baseline definition, but applies to a broader class of sustainable strategic relationships.
7 Conclusion
With a non-exiting but strategically responsive exploitee, SEE refines MPE by imposing viability and renegotiation-proof exploiter-optimal selection. In the hegemon–client example, SEE yields restrained extraction that keeps the regime just at the viability boundary. More generally, the framework applies whenever one party cannot exit but retains an effort margin, including colonial settings, labor relationships and locked-in suppliers. By combining dynamic sustainability with renegotiation-proofness, SEE offers a compact refinement of MPE for analyzing persistent exploitative interactions in economics and political economy.