### v2

# Sustainable Exploitation Equilibria for Dynamic Games

Abstract
We introduce the Sustainable Exploitation Equilibrium (SEE), a refinement of Markov Perfect Equilibrium (MPE) for dynamic games with an exploiter–exploitee structure. SEE imposes renegotiation-proof exploiter-optimal selection on the set of rationally viable stationary Markov equilibria, where viability follows from sequential rationality when exiting a sustainability set entails catastrophic losses. Unlike MPE, SEE rules out equilibria in which the exploiter optimally drives the state to collapse despite positive continuation payoffs. The exploitee cannot exit, but retains a strategic effort margin affecting dynamics and payoffs. We establish existence under standard conditions, and the refinement is illustrated in a hegemon–client model of foreign politics.
Keywords: Markov perfect equilibrium, equilibrium refinements, viability, renegotiation-proofness, sustainability
JEL: C72, C73, D74.
1 Introduction
Exploitation arises when one agent extracts value from another whose continued functioning is essential for the relationship itself. In such dynamic environments, standard equilibrium concepts such as Markov Perfect Equilibrium (MPE) may admit outcomes in which the exploiter optimally extracts too aggressively, driving the state into collapse despite positive continuation payoffs. These equilibria rely on the exploiter being indifferent to catastrophic breakdown once it occurs. To address this, this paper introduces the Sustainable Exploitation Equilibrium (SEE), a refinement of stationary Markov Perfect Equilibrium that rules out such collapse outcomes by deriving viability from rationality. When exiting a sustainability region entails sufficiently large losses for the exploiter—and when a feasible safe action exists that preserves the relationship—any equilibrium placing positive probability on collapse is sequentially irrational. Under these conditions, restricting attention to viable stationary equilibria is without loss.
SEE then further refines the set of rationally viable stationary MPE by imposing renegotiation-proofness and exploiter-optimal selection. SEE is motivated by settings where one party cannot exit but remains responsive, such as in hegemon–client dynamics, as the exploitee retains a strategic effort margin that shapes both payoffs and dynamics. In this way, SEE formalizes the idea of “sustainable exploitation”: the exploiter extracts as much as possible while ensuring the long-run survival and productivity of the exploitee. This rational-viability foundation distinguishes SEE from equilibrium selection rules that impose sustainability as an external constraint.
2 Related Literature
Our work connects three strands of existing literature.
Equilibrium refinements.
SEE builds on the refinement tradition in dynamic games. (MaskinTirole2001a) develop the theory of Markov perfect equilibrium; (Bernheim1987) introduce coalition-proof Nash equilibrium; (Selten1975) trembling-hand perfection; and (FarrellMaskin1989) renegotiation-proofness in repeated games.
Viability and dynamic control.
The viability constraint draws from (Aubin1991) Viability Theory and constrained Markov decision processes (Altman1999), ensuring state trajectories remain within sustainability sets. This connects to the resource economics literature, e.g. (Dasgupta1979), (Chichilnisky1997), and (Dockner2000), where long-run viability is central to dynamic exploitation of resources.
Political economy and relational contracts.
SEE also relates to political economy models of extraction and state capacity ((AcemogluRobinson2006); (BesleyPersson2011)) and to relational contracting without full commitment ((ThomasWorrall1988); (Levin2003)). In these contexts, weaker parties cannot exit but retain effort margins that shape dynamics. SEE formalizes the logic of “restrained exploitation” within a Markovian framework.
3 Model and Definition
Environment and timing.
Time is discrete, $t=0,1,2,\dots$. The state $s_{t}\in S$ evolves on a compact metric space $S$. In each period:
  1. 1.
    The exploiter $X$ observes $s_{t}$ and chooses extraction $x_{t}\in X(s_{t})$.
  2. 2.
    The exploitee $E$ observes $(s_{t},x_{t})$ and chooses effort $e_{t}\in E(s_{t})$.
  3. 3.
    The state transitions to $s_{t+1}\sim Q(\cdot\mid s_{t},x_{t},e_{t})$ (deterministic dynamics are $s_{t+1}=f(s_{t},x_{t},e_{t})$).
Per-period payoffs are $u^{X}(s_{t},x_{t},e_{t})$ and $u^{E}(s_{t},x_{t},e_{t})$, bounded and continuous. Both players discount with $\delta\in(0,1)$.
Roles of the players
Definition 1 (Exploiter).
The exploiter $X$ directly extracts value via $x\in X(s)$, appropriating resources and typically reducing the future state. $X$’s stage payoff $u^{X}(s,x,e)$ is weakly increasing in $x$ for fixed $(s,e)$. A Markov policy for $X$ is $\sigma^{X}:S\to X(s)$.
Definition 2 (Exploitee (no exit, effort choice)).
The exploitee $E$ cannot exit the relationship (no outside option) but chooses effort $e\in E(s)$ after observing $(s,x)$. Effort affects contemporaneous payoffs and/or the law of motion. A Markov policy for $E$ is $\sigma^{E}:S\times X(s)\to E(s)$ (second-mover response).
Equilibrium notion (Markov–Stackelberg Equilibrium).
A Markov–Stackelberg Equilibrium (MSE) consists of policies $(\sigma^{X},\sigma^{E})$ such that: (i) for each $(s,x)$, $\sigma^{E}(s,x)$ is a best response to $\sigma^{X}$ given follower timing; (ii) $\sigma^{X}(s)$ anticipates $\sigma^{E}$ and is a best response given $E$’s continuation behavior. Values $(W^{X},W^{E})$ solve the associated Bellman system.
Rational interpretation of viability.
We interpret transitions outside a sustainability set $V\subset S$ as catastrophic breakdown for the exploiter, in the sense that sufficiently small reductions in extraction can strictly improve continuation payoffs by preserving the relationship. Under this interpretation, equilibrium behavior that places positive probability on exiting $V$ is incompatible with sequential rationality when collapse entails sufficiently large losses.
Definition 3 (Sustainable Exploitation Equilibrium (SEE)).
Let $V\subseteq S$ be a nonempty compact viability set that is invariant under admissible state transitions. Let $\mathcal{E}$ denote the set of all stationary Markov Perfect Equilibria, and let $\mathcal{E}_{V}\subseteq\mathcal{E}$ denote the subset whose induced state process remains in $V$. Under rational viability (formalized below), restricting attention to $\mathcal{E}_{V}$ is without loss.
A Sustainable Exploitation Equilibrium is any stationary MPE $\sigma^{*}\in\mathcal{E}$ satisfying:
  1. (i)
    Viability: The state induced by $\sigma^{*}$ remains in the viability set $V$ for all $s\in V$ (i.e., $\sigma^{*}\in\mathcal{E}_{V}$).
  2. (ii)
    Renegotiation-proofness: At any state $s\in V$, there is no alternative viable profile $\tilde{\sigma}\in\mathcal{E}_{V}$ such that:
    $W_{E}(\tilde{\sigma};s)\geq W_{E}(\sigma^{*};s)\quad\text{and}\quad W_{X}(\tilde{\sigma};s)\geq W_{X}(\sigma^{*};s),$
    with at least one inequality strict.
  3. (iii)
    Exploiter-optimality: Among all renegotiation-proof elements of $\mathcal{E}_{V}$, $\sigma^{*}$ maximizes the exploiter’s value $W_{X}(\cdot;s)$ at the state $s$ (with ties broken lexicographically).
Any profile satisfying (i)–(iii) is a Sustainable Exploitation Equilibrium.
4 Relation to Nash and Markov Perfect Equilibrium
We position the Sustainable Exploitation Equilibrium (SEE) within the standard equilibrium hierarchy.
Nash equilibrium (NE).
A strategy profile $\sigma=(\sigma_{X},\sigma_{E})$ is a Nash equilibrium if no player can profitably deviate unilaterally. NE is broad and may rely on history-dependent strategies.
Subgame perfect equilibrium (SPE).
A refinement of NE, SPE requires sequential rationality in every subgame. This eliminates non-credible threats. Clearly,
$\text{SPE}\subseteq\text{NE}.$
Markov perfect equilibrium (MPE).
A further refinement of SPE for dynamic games with a payoff-relevant state $s\in S$. In MPE, strategies are Markovian: $\sigma_{i}(s)$ depends only on the current state, not the full history. For two-player sequential-move environments, the relevant benchmark is a Markov–Stackelberg equilibrium (MSE): the follower chooses $\sigma_{E}(s,x)$ as a best response given the state and leader’s action, while the leader anticipates this response in selecting $\sigma_{X}(s)$. Thus,
$\text{MPE (MSE)}\subseteq\text{SPE}\subseteq\text{NE}.$
Sustainable Exploitation Equilibrium (SEE).
SEE is defined as a viable, renegotiation-proof MPE with exploiter-optimal selection. Formally, a stationary Markov profile $(\sigma_{X},\sigma_{E})$ is an SEE if the aforementioned conditions of Definition 3 are satisfied.
It follows immediately that
$\text{SEE}\subseteq\text{MPE}\subseteq\text{SPE}\subseteq\text{NE}.$
5 Existence under rational viability
We derive viability from rationality by interpreting exits from a sustainability set as catastrophic for the exploiter. We interpret the catastrophe penalty $M$ as a reduced-form representation of franchise loss, regime collapse, or irreversible destruction of the exploitee following breakdown of the relationship.
Assumption (Catastrophic collapse and safe action). Payoffs are bounded: $|u^{X}(s,x,e)|\leq\bar{u}<\infty$. Let $V\subset S$ be a nonempty compact set. For each $s\in V$, there exists a safe action $\underline{x}(s)\in X(s)$ such that, under the exploitee’s best response,
$Q\!\left(V\mid s,\underline{x}(s),\sigma_{E}(s,\underline{x}(s))\right)=1.$
Theorem 1 (Existence of SEE under rational viability).
Suppose the state space $S$ and action sets $X(s),E(s)$ are compact, payoffs are bounded and continuous, and the transition kernel $Q$ satisfies the Feller property. Interpret $S\setminus V$ as a collapse region generating a catastrophic loss for the exploiter.
Then there exists $M^{\star}<\infty$ such that, for all catastrophe penalties $M\geq M^{\star}$, every stationary Markov Perfect Equilibrium of the penalized game
$\hat{u}^{X}(s,x,e)\equiv u^{X}(s,x,e)-M\cdot\mathbf{1}\{s^{\prime}\notin V\},\qquad s^{\prime}\sim Q(\cdot\mid s,x,e),$
is viable. Moreover, among viable stationary Markov Perfect Equilibria, a nonempty compact subset of renegotiation-proof equilibria exists, and an exploiter-optimal selection exists. Hence a Sustainable Exploitation Equilibrium exists.
Sketch.
Under compactness and continuity, stationary Markov Perfect Equilibria exist. Fix any $s\in V$. If an equilibrium action induces $Q(S\setminus V\mid s,x,e)>0$, the exploiter’s expected one-period payoff is at most $\bar{u}-Mp$ for some $p>0$. Under the safe action $\underline{x}(s)$, the expected penalty is zero and the payoff is at least $\bar{u}$. For $M$ sufficiently large, the safe action strictly dominates any action with positive exit probability. By continuity and compactness, a finite threshold $M^{\star}$ exists. Hence any equilibrium must be viable. Existence of renegotiation-proof and exploiter-optimal selections follows by compactness and continuity. ∎
6 Example: Foreign Politics (Hegemon–Client SEE)
6.1 Primitives and timing
Two players: a hegemon $H$ (the exploiter) and a client state $C$ (the exploitee). The state $s_{t}\in S\subset\mathbb{R}_{+}$ represents client political capacity/stability (e.g., tax capacity, administrative control, consent). Time is discrete $t=0,1,2,\dots$.
Within each period:
  1. 1.
    $H$ observes $s_{t}$ and chooses an extraction/pressure level $x_{t}\in X(s_{t})$ (e.g., tribute, resource concessions, policy compliance).
  2. 2.
    $C$ observes $(s_{t},x_{t})$ and chooses effort $e_{t}\in E(s_{t})$ (e.g., governance effort, compliance/implementation, regime maintenance).
  3. 3.
    The state updates to $s_{t+1}=f(s_{t},e_{t})-h(x_{t})$, where $f$ is increasing in $s$ and $e$ and $h$ is increasing in $x$; or more generally $s_{t+1}\sim Q(\cdot\mid s_{t},x_{t},e_{t})$.
Both discount with factor $\delta\in(0,1)$. The client cannot exit the relationship (no outside option) but has strategic effort. A viability requirement $s_{t}\geq s_{\min}$ avoids regime collapse or state failure.
6.2 Payoffs and feasibility
Per-period payoffs are continuous and bounded:
$u^{H}(s,x,e)\;=\;\pi(x)\;-\;k(s,e),\qquad u^{C}(s,x,e)\;=\;b(s)\;-\;\phi(e)\;-\;d(x),$
with $\pi^{\prime}(x)>0$, $k\geq 0$ (costs to monitor/coerce or instability spillovers), $b^{\prime}(s)\geq 0$, $\phi$ convex with $\phi^{\prime}(0)=0$, and $d^{\prime}(x)\geq 0$. Feasibility (viability) imposes
$x\;\leq\;f(s,e)-s_{\min}.$
6.3 Equilibrium notion and SEE
We adopt a Markov–Stackelberg Equilibrium (MSE): the follower $C$ best responds with $\sigma^{C}(s,x)$, and the leader $H$ best responds with $\sigma^{H}(s)$ anticipating $\sigma^{C}$. A Sustainable Exploitation Equilibrium (SEE) is a viable, renegotiation-proof MSE with exploiter-optimal selection (as in Definition 1).
6.4 Follower’s best response
Given $(s,x)$ and anticipating future play, $C$ solves
$\max_{e\in E(s)}\;\Big\{b(s)-\phi(e)-d(x)+\delta W^{C}\!\big(f(s,e)-h(x)\big)\Big\}.$
An interior first-order condition is
$\phi^{\prime}(e^{*})\;+\;\delta\,W^{C}_{s}\!\big(s^{\prime}\big)\,f_{e}(s,e^{*})\;=\;0,\qquad s^{\prime}\equiv f(s,e^{*})-h(x).$
Under standard concavity/regularity, the implicit function theorem yields $\partial e^{*}/\partial x$.
6.5 Leader’s SEE problem
Anticipating $e^{*}(s,x)$, the hegemon solves
$V^{H}(s)\;=\;\max_{x\in X(s)}\;\Big\{\pi(x)-k(s,e^{*}(s,x))+\delta\,V^{H}\!\big(f(s,e^{*}(s,x))-h(x)\big)\Big\}$
subject to viability $x\leq f(s,e^{*}(s,x))-s_{\min}$. Let $\mu\geq 0$ be the multiplier on viability. The Lagrangian is
$\mathcal{L}(x;s,\mu)\;=\;\pi(x)-k(s,e^{*})+\delta V^{H}(s^{\prime})\;+\;\mu\big(f(s,e^{*})-h(x)-s_{\min}\big),$
where $e^{*}\equiv e^{*}(s,x)$ and $s^{\prime}\equiv f(s,e^{*})-h(x)$.
The (interior) first-order condition with envelope terms from $e^{*}(s,x)$ is
$\pi^{\prime}(x)\;-\;\delta V^{H}_{s}(s^{\prime})\,\big(h^{\prime}(x)-f_{e}(s,e^{*})\,\partial e^{*}/\partial x\big)\;-\;\mu\,\big(h^{\prime}(x)-f_{e}(s,e^{*})\,\partial e^{*}/\partial x\big)\;=\;0,$
with complementary slackness $\mu\big(f(s,e^{*})-h(x)-s_{\min}\big)=0$.
Interpretation.
The hegemon equates marginal extraction benefits $\pi^{\prime}(x)$ to the discounted shadow cost of lost future capacity $h^{\prime}(x)$ net of the induced effort response $f_{e}\partial e^{*}/\partial x$. When viability binds, the shadow price $\mu>0$ further restrains extraction.
6.6 Stationary SEE and regimes
A stationary SEE consists of $(x^{*}(s),e^{*}(s,x))$ and an invariant distribution supported on $V=\{s\geq s_{\min}\}$. Under rational viability, two canonical regimes arise:
  • Interior (viability slack, $\mu=0$). If the unconstrained fixed point $s^{*}$ of $s^{\prime}=f(s,e^{*}(s,x^{*}))-h(x^{*})$ satisfies $s^{*}>s_{\min}$, the FOCs imply a generalized “golden rule” where effective regeneration $f_{s}+f_{e}\,\partial e^{*}/\partial s$ balances discounting (details depend on primitives).
  • Boundary (viability binds, $\mu>0$). If the unconstrained $s^{*}$ would fall below $s_{\min}$, SEE selects $s^{*}=s_{\min}$ and $x^{*}=f(s_{\min},e^{*}(s_{\min},x^{*}))-s_{\min}$, with $\mu$ set to satisfy the FOC.
Additional Remark on Outside Option
One might alternatively assume that the exploitee $E$ has an outside option $U^{out}_{E}(s)$, in which case viability must be supplemented by an individual rationality (IR) constraint: $W^{E}(s)\geq U^{out}_{E}(s)$ for all $s\in V$. This yields an even tighter refinement, as it restricts feasible equilibria to those in which the exploitee weakly prefers participation over exit.
7 Conclusion
With a non-exiting but strategically responsive exploitee, SEE refines Markov Perfect Equilibrium by deriving viability from rationality and imposing renegotiation-proof exploiter-optimal selection. Unlike MPE, SEE eliminates equilibria in which the exploiter optimally drives the state to collapse despite positive continuation payoffs. In equilibrium, SEE yields either an interior “golden-rule”–type extraction path or, when unconstrained extraction would cause collapse, a boundary outcome in which the state is pinned at the viability threshold. The framework applies broadly to settings in which one party cannot exit but retains an effort margin—such as colonial relationships, labor arrangements, and locked-in suppliers—providing a compact refinement for analyzing persistent exploitative interactions in economics and political economy.